OpenStudy (adi3):

Will Medal!!! Please help Suppose f(x) = 3x+1 and g(x) = x^2 +2x. Find x such that f(g(x)) = 10.

OpenStudy (adi3):

@ganeshie8

OpenStudy (anonymous):

Use Wolfram alpha for this

OpenStudy (adi3):

wolfram alpha is not correct

OpenStudy (danjs):

replace each x value in f(x) with the entire function g(x)

OpenStudy (danjs):

f(x) = 3x+1 f(g(x)) = 3*g(x) + 1

OpenStudy (adi3):

What do you mean by that?

OpenStudy (danjs):

g(x) = x^2+2x f(g(x)) = 3*g(x) + 1 = 3 * (x^2 + 2x) + 1

OpenStudy (danjs):

see that

OpenStudy (adi3):

yeah i get it

OpenStudy (danjs):

then they tell you that equals to a value and find x

OpenStudy (adi3):

it equals 3x^2 + 6x + 1 right

OpenStudy (danjs):

10 = 3x^2 + 6x + 1

OpenStudy (danjs):

yeah the f(g(x)) does, then they tell you that is equal to 10

OpenStudy (adi3):

then how do you solve it after that, that s the part which is the most confusing for me

OpenStudy (danjs):

it is a quadratic, to get the solution, you put it in standard form first.. ax^2+bx+c=0

OpenStudy (danjs):

3x^2 + 6x - 9 = 0 3*( x^2 + 2x - 3) = 0

OpenStudy (adi3):

then what to do next

OpenStudy (danjs):

so, x^2 + 2x -3 = 0 you can always just apply the quadratic formula, but this one looks to factor easy

OpenStudy (danjs):

(x - 1)*(x+3) = 0

OpenStudy (adi3):

wait from where did x-1 come from

OpenStudy (danjs):

so, that is true if.. x - 1 = 0 OR x + 3 = 0

OpenStudy (danjs):

x = 1 or x=-3

OpenStudy (danjs):

factoring a quadratic

OpenStudy (adi3):

ohh thanks

OpenStudy (danjs):

distribute the thing and check , it goes back to the original

OpenStudy (danjs):

or simply just plug in the coefficients to the quadratic equation and hit a button to solve for you

OpenStudy (adi3):

i did not get the part after this 3x^2 + 6x - 9 = 0 3*( x^2 + 2x - 3) = 0

OpenStudy (adi3):

what did you do after this

OpenStudy (danjs):

divide both sides with a 3

OpenStudy (danjs):

0/3 = 0

OpenStudy (danjs):

Once you get down to standard form ax^2 + bx + c = 0 solving that expression for x will give you the quadratic formula

OpenStudy (adi3):

yeah i get it, thanxs

OpenStudy (danjs):

\[x = \frac{ -b \pm \sqrt{b^2 - 4*a*c} }{ 2a }\]

OpenStudy (danjs):

k, welcome

OpenStudy (adi3):

thnxs